• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name M. Cummings

    Log entry time 08:05:12 on March15,2012

    Entry number 366543

    keyword=Owl Shift Summary

    Owl Shift Summary

    Shift Crew:
    SL: Melissa Cummings
    TO: Dustin Keller
    3rd: Hyekoo Kang

    00:15 Start of shift: MCC reports FSD / software issues, in contact with software person, no estimated time for beam at present
    00:30 MCC sends tune beam
    01:00 Hall C changed IHWP to OUT
    03:00 MCC still tuning beam...
    04:00 Hall C changed IHWP2 (passive) to IN, IHWP is still OUT.
    07:20 Hall C changed IHWP to IN, IHWP2 (passive) changed to OUT
    08:00 End of shift, still waiting for MCC to finish tuning beam.